# Actions Speak Louder Than Words: Rate-Reward Trade-off in Markov Decision Processes

### v2

@@ 3 Preliminaries and System Model
[…]
@@
Let $\mathcal{M}$ denote the set of messages, with each message $m$ uniformly sampled from $\mathcal{M}$. The encoder is defined as a sequence of mappings, $\mathcal{E}\triangleq\{\mathcal{E}_{1},\ldots,\mathcal{E}_{n}\}$, where each mapping $\mathcal{E}_{t}:\mathcal{M}\times{\mathcal{S}}^{t}\times\mathcal{X}^{t-1}\times{\mathcal{{Y}}}^{t-1}\to{\mathcal{X}}$ generates the channel input at time $t$. In other words, the channel input at time $t$, $x_{t}=\mathcal{E}_{t}({m},\bm{s^{t}},\bm{x^{t-1}},\bm{{y}^{t-1}})$, is a function of $m$ and all the historical information available at the transmitter up to time $t$. The decoder is defined as the mapping, $\mathcal{D}:\mathcal{Y}^{n}\to{\mathcal{M}}$, which reconstructs the message from all $n$ channel outputs, $\hat{m}=\mathcal{D}(\bm{y}^{n})$. The pair $(\mathcal{E},\mathcal{D})$ constitutes a code, where $n$ is called the code length. Suppose the message set is =$\mathcal{M}=\{1,2,3,\ldots,2^{k}\}$, then each message can be represented with $k$ bits. The rate of the code $(\mathcal{E},\mathcal{D})$ is defined as $R_{\mathcal{E},\mathcal{D}}=k/n$.
@@
$\vspace{-3pt}\begin{array}[]{cl}\textbf{P2:}&\max\limits_{{\mathcal{E},\mathcal{D}}}R_{\mathcal{E},\mathcal{D}},\quad\text{subject to }P_{e}^{(n)}\leq\sigma.\end{array}\vspace{-3pt}$ (3)
@@
(b) Controller & Encoder: Within this framework, the MDP controller and the FSC encoder represent the two aspects of the same entity, jointly responsible for selecting an action $x_{t}\in\mathcal{X}$ at each time step. However, their objectives differ: the controller aims to maximize the reward, while the encoder seeks to maximize the message rate. Unlike the controller, which can focus on stationary deterministic policies, the encoder must account for more complex policy forms. For any message $m$, the encoder described in the previous part can be viewed as a history-dependent policy for the MDP. Therefore, we consider the joint control and coding policy in its most general form. The policy $\mathcal{E}$ is represented as a sequence of mappings $\{\mathcal{E}_{i}:0\leq i\leq n-1\}$, where $\mathcal{E}_{i}$ is defined as $\mathcal{E}_{i}:\mathcal{M}\times{\mathcal{S}}^{i}\times\mathcal{X}^{i-1}\to{\mathcal{X}}$. Each message is transmitted via a sequence of $n$ actions. For example, if the controller begins to transmit a message $m$ at time $t$, then $x_{t+i}=\mathcal{E}_{i}(m,\bm{s}_{t}^{t+i},\bm{x}_{t}^{t+i-1})$ for $10\leq i\leq n-$ 1. We assume the controller always has a new message ready for transmission immediately after completing the transmission of the previous message, and each message is uniformly sampled from $\mathcal{M}$. The long-term average reward of the MDP under policy $\mathcal{E}$ is denoted by $G_{\mathcal{E}}$.
@@
[…]
$s.t.$ $\ \ G_{\mathcal{E}}\geq V\quad\text{and}\quad P_{e}^{(n)}\leq\sigma.\vspace{-3pt}$ (6)
@@ 4 The Capacity-Reward Trade-off
$\vspace{-3pt}I(X;S^{+}|S)=\mathbb{E}_{p(x,s^{+},s)}\left[\log\frac{p(s,s^{+}|s)}{p(x|s)p(s^{+}|s)}\right].\vspace{-3pt}$ (7)
@@
$\max_{w\in\mathcal{W}}$ $I(w,\bm{T})$
$s.t.$ $\sum_{s\in\mathcal{S}}\sum_{x\in\mathcal{X}}w(s,x)r(s,x)\geq V$
@@
$I(w,\bm{T})\triangleq\sum_{s\in\mathcal{S}}\sum_{x\in\mathcal{X}}w(s,x)\sum_{s^{\prime}\in\mathcal{S}}\bm{T}(s^{\prime}|s,x)\log\frac{\bm{T}(s^{\prime}|s,x)\sum_{x^{\prime\prime}}w(s,x^{\prime\prime})}{\sum_{x^{\prime}}\bm{T}(s^{\prime}|s,x^{\prime})w(s,x^{\prime})}.\vspace{-5pt}$
@@ 5 Act2Comm: A practical coding scheme
This section presents Act2Comm, a learning-based practical coding scheme that balances both control and communication objectives. This framework assumes a pre-determined control policy $\pi$ that satisfies the reward constraint $G_{\pi}\geq V$, referred to as the target policy. Such a target policy can be easily derived using traditional MDP or RL algorithms. Alternatively, solving the problem in Theorem 2 yields a policy $\pi$ that precisely satisfies $G_{\pi}=V$. Act2Comm aims to learn a coding policy that: (1) closely mimics the stochastic behavior of the target policy; (2) minimizes the probability of decoding errors for a given coding rate and a finite code length. That is, Act2Comm takes a policy achieving the desired reward, and embeds messages into it with the desired reliability. For ease of reference, we denote the element of matrix $\bm{X}$ at the $i$-th row and $j$-th column as ${X}[i,j]$. The sets of states and actions are indexed as $\mathcal{S}=\{0,1,\ldots,|\mathcal{S}|-1\}$ and $\mathcal{X}=\{0,1,\ldots,|\mathcal{X}|-1\}$.
@@
$\bm{U}=\mathcal{Q}(|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})),\vspace{-2pt}$ (12)
@@
Let $\pi$ denote the target policy, with $\pi(x|s)$ representing the probability of taking action $x$ in state $s$. Let $f_{U}(x|s)$ denote the frequency of selecting $x$ in state $s$ across all decision rules in $\bm{U}$. We then use the mean square error (MSE) between $\pi$ and $f_{U}$ to measure the control loss for its stability in experiments. However, $f_{U}$ is non-differentiable during the backpropagation as it is discrete. To address this issue, we estimate $f_{U}(x|s)$ using $\bm{Z}$ in equation 12. Let $\bm{e}$ denote the all-one row vector, and define $\Gamma_{Z}(\bm{T},s,x)\triangleq\text{Sigmoid}(\gamma(|\mathcal{X}|\text{Sigmoid}(\bm{Z}[s,:])-x\bm{e}))$. When $\gamma>0$ is sufficiently large, $\Gamma_{Z}(\bm{T},s,x)$ is a $(kR)$-dim vector with elements close to either 00 or $1$. Additionally, $\Gamma_{Z}(\bm{T},s,x)\bm{e}^{\top}$ approximates the number of elements in $\bm{U}[s,:]$ that are not less than $x$. We refer to $\gamma$ as the temperature parameter and estimate $f_{U}(x|s)$ for $x>0$ as follows:
$\vspace{-2pt}f_{U}(x|s)\approx\hat{f}_{U}(x|s)\triangleq\frac{1}{kR}\left[\Gamma_{Z}(\bm{T},s,x-1)\bm{e}^{\top}-\Gamma_{Z}(\bm{T},s,x)\bm{e}^{\top}\right].\vspace{-2pt}$ (13)
For $x=0$, we have $f_{U}(0|s)\approx\hat{f}_{U}(0|s)\triangleq 1-\Gamma_{Z}(\bm{T},s,0)\bm{e}^{\top}/kR$. As a result, we define the control loss as $\mathcal{L}_{cont}=\textit{MSE}(\pi,\hat{f})$.
@@ Appendix A Notation and Definitions
${S^{+}}$, ${S}$ Future and current state (random variables)
@@
${S^{+}}$, ${S}$ Future and current state (random variables)
@@
${S^{+}}$, ${S}$ Future and current state (random variables)
@@ B.1 Proof of Theorem 1
=$\overset{(a)}{=}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\left[H(S_{i+1}|S^{i}_{2},s_{1})-H(S_{i+1}|S^{i}_{2},U^{N},X_{i},s_{1})\right]$
$=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\left[H(S_{i+1}|S_{i},s_{1})-H(S_{i+1}|S_{i},X_{i},s_{1})\right]$
@@
where (a) holds because $X_{i}$ is determined by $S^{i}$ and $U_{i}$. It follows that
@@
=$\overset{(a)}{=}\max_{\{p(x_{i}|s^{i},x^{i-1})\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1})$
=$\overset{(b)}{=}\max_{\{p(x_{i}|s_{i})\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1})$ (15)
In equation B.1, (a) follows from the equivalence between the action-state channel and the EAS channel. We next prove (b). Let $\eta$ denote a history-dependent encoder and $\{\eta_{i}(x_{i}|s^{i},x^{i-1})\}_{i\geq 1}$ denote the associated input distribution. Furthermore, let $P^{\eta}_{i}(s,x)$ denote the probability that $S_{i}=s$ and $X_{i}=s$ conditioned on the encoder $\eta$. Note that $\eta$ can be viewed as a history-dependent policy for the MDP. Then according to MDP theory, for any history-dependent policy $\eta$, there exists a Markov policy $\eta^{\prime}$ such that $\eta$ and $\eta^{\prime}$ share the same joint probability distribution of states and actions. In particular, denote by $\eta^{\prime}_{i}(x_{i}|s_{i})$ the probability of selecting action $x_{i}$ given that the state is $s_{i}$ at time $i$. Then $\eta^{\prime}$ can be seen as a Markov encoder with input distribution $\{\eta^{\prime}_{i}(x_{i}|s_{i})\}_{i\geq 1}$. By letting
@@ C.3 Training and inference algorithm
32:        $\mathcal{L}_{cont}=\text{MSE}(\bm{\pi},\bm{\hat{f}})$ // Control loss term for the estimated $\hat{f}$
@@
42:     $\mathcal{L}_{com}=\text{cross-entropy}(\bm{m},\bm{\hat{M}})$ // Communication loss term
@@
2:  $\mathbf{C^{(1)}}\in 0^{1\times\frac{\mu}{R}\times 3}$, // Initialize the feedback block
@@
14:  $\bm{Z}=[\bm{z^{(1)}};\ldots;\bm{z^{(l)}}]\in\mathbb{R}^{\frac{k}{\mu}\times\frac{\mu|S|}{R}}\rightarrow\mathbb{R}^{|S|\times\frac{k}{R}}$, // Collect the final codeword
@@
16:  $\bm{s}=[s_{1};\ldots;s_{\frac{k}{R}}]\in\mathcal{S}^{\frac{k}{R}}\rightarrow\mathcal{S}^{\frac{k}{\mu}\times\frac{\mu}{R}}$, // Collect the observed states
17:  $\bm{\hat{M}}=\mathcal{D}(\bm{s})\in\mathbb{R}^{\frac{k}{\mu}\times 2^{\mu}}$, // Decode the logits
18:  $\bm{\hat{m}}=\text{argmax}(\text{Softmax}(\bm{\hat{M}}),\text{dim}=-1)$ // Decode each block labels
@@
[…]
Table 4: Number of parameters, FLOPs, and coding times for the Act2Comm scheme ($K=12$, $\mu=3$, $R=1/3$, $s_{in}=20$) with varying critic network training steps $s_{in}$ for both training and inference phases. Note that the backpropagation computational complexity is approximately 23232–32 – 3 times that of the forward computation.

### v3

@@ 3 Preliminaries and System Model
$\textbf{P1:}\quad\max_{\pi}\ \lim_{N\rightarrow\infty}\frac{1}{N}\mathbb{E}\left[\sum_{t=1}^{N}r(s_{t},x_{t})|s_{1}\sim\alpha\right].$ (1)
@@
Let $\mathcal{M}$ denote the set of messages, with each message $m$ uniformly sampled from $\mathcal{M}$. The encoder is defined as a sequence of mappings, $\mathcal{E}\triangleq\{\mathcal{E}_{1},\ldots,\mathcal{E}_{n}\}$, where each mapping $\mathcal{E}_{t}:\mathcal{M}\times{\mathcal{S}}^{t}\times\mathcal{X}^{t-1}\times{\mathcal{{Y}}}^{t-1}\to{\mathcal{X}}$ generates the channel input at time $t$. In other words, the channel input at time $t$, $x_{t}=\mathcal{E}_{t}({m},\bm{s^{t}},\bm{x^{t-1}},\bm{{y}^{t-1}})$, is a function of $m$ and all the historical information available at the transmitter up to time $t$. The decoder is defined as the mapping, $\mathcal{D}:\mathcal{Y}^{n}\to{\mathcal{M}}$, which reconstructs the message from all $n$ channel outputs, $\hat{m}=\mathcal{D}(\bm{y}^{n})$. The pair $(\mathcal{E},\mathcal{D})$ constitutes a code, where $n$ is called the code length. Suppose the message set is $\mathcal{M}=\{1,2,3,\ldots,2^{k}\}$, then each message can be represented with $k$ bits. The rate of the code $(\mathcal{E},\mathcal{D})$ is defined as $R_{\mathcal{E},\mathcal{D}}=k/n$.
@@
$\vskip-3.0pt\begin{array}[]{cl}\textbf{P2:}&\max\limits_{{\mathcal{E},\mathcal{D}}}R_{\mathcal{E},\mathcal{D}},\quad\text{subject to }P_{e}^{(n)}\leq\sigma.\end{array}\vskip-3.0pt$ (3)
@@
(b) Controller & Encoder: Within this framework, the MDP controller and the FSC encoder represent the two aspects of the same entity, jointly responsible for selecting an action $x_{t}\in\mathcal{X}$ at each time step. However, their objectives differ: the controller aims to maximize the reward, while the encoder seeks to maximize the message rate. Unlike the controller, which can focus on stationary deterministic policies, the encoder must account for more complex policy forms. For any message $m$, the encoder described in the previous part can be viewed as a history-dependent policy for the MDP. Therefore, we consider the joint control and coding policy in its most general form. The policy $\mathcal{E}$ is represented as a sequence of mappings $\{\mathcal{E}_{i}:0\leq i\leq n-1\}$, where $\mathcal{E}_{i}$ is defined as $\mathcal{E}_{i}:\mathcal{M}\times{\mathcal{S}}^{i}\times\mathcal{X}^{i-1}\to{\mathcal{X}}$. Each message is transmitted via a sequence of $n$ actions. For example, if the controller begins to transmit a message $m$ at time $t$, then $x_{t+i}=\mathcal{E}_{i}(m,\bm{s}_{t}^{t+i},\bm{x}_{t}^{t+i-1})$ for $0\leq i\leq n-1$. We assume the controller always has a new message ready for transmission immediately after completing the transmission of the previous message, and each message is uniformly sampled from $\mathcal{M}$. The long-term average reward of the MDP under policy $\mathcal{E}$ is denoted by $G_{\mathcal{E}}$.
@@
$\textbf{P3:}\quad\max\limits_{\mathcal{E,D}}$ $\ \ R_{\mathcal{E,D}}$ (5)
$s.t.$ $\ \ G_{\mathcal{E}}\geq V\quad\text{and}\quad P_{e}^{(n)}\leq\sigma.\vskip-3.0pt$ (6)
@@ 4 The Capacity-Reward Trade-off
$\vskip-3.0ptI(X;S^{+}|S)=\mathbb{E}_{p(x,s^{+},s)}\left[\log\frac{p(s,s^{+}|s)}{p(x|s)p(s^{+}|s)}\right].\vskip-3.0pt$ (7)
@@
$\max_{w\in\mathcal{W}}\$$I(w,\bm{T})$
$s.t.\$$\sum_{s\in\mathcal{S}}\sum_{x\in\mathcal{X}}w(s,x)r(s,x)\geq V$
@@
$I(w,\bm{T})\triangleq\sum_{s\in\mathcal{S}}\sum_{x\in\mathcal{X}}w(s,x)\sum_{s^{\prime}\in\mathcal{S}}\bm{T}(s^{\prime}|s,x)\log\frac{\bm{T}(s^{\prime}|s,x)\sum_{x^{\prime\prime}}w(s,x^{\prime\prime})}{\sum_{x^{\prime}}\bm{T}(s^{\prime}|s,x^{\prime})w(s,x^{\prime})}.\vskip-5.0pt$
@@ 5 Act2Comm: A practical coding scheme
This section presents Act2Comm, a learning-based practical coding scheme that balances both control and communication objectives. This framework assumes a pre-determined control policy $\pi$ that satisfies the reward constraint $G_{\pi}\geq V$, referred to as the target policy. Such a target policy can be easily derived using traditional MDP or RL algorithms. Alternatively, solving the problem in Theorem 2 yields a policy $\pi$ that precisely satisfies $G_{\pi}=V$. Act2Comm aims to learn a coding policy that: (1) closely mimics the stochastic behavior of the target policy; (2) minimizes the probability of decoding errors for a given coding rate and a finite code length. That is, Act2Comm takes a policy achieving the desired reward, and embeds messages into it with the desired reliability. For ease of reference, we denote the element of matrix $\bm{X}$ at the $i$-th row and $j$-th column as X[i,j]{X}[i,j]. The sets of states and actions are indexed as $\mathcal{S}=\{0,1,\ldots,|\mathcal{S}|-1\}$ and $\mathcal{X}=\{0,1,\ldots,|\mathcal{X}|-1\}$.
@@
$\bm{U}=\mathcal{Q}(|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})),\vskip-2.0pt$ (12)
@@
Let $\pi$ denote the target policy, with $\pi(x|s)$ representing the probability of taking action $x$ in state $s$. Let $f_{U}(x|s)$ denote the frequency of selecting $x$ in state $s$ across all decision rules in $\bm{U}$. We then use the mean square error (MSE) between $\pi$ and $f_{U}$ to measure the control loss for its stability in experiments. However, $f_{U}$ is non-differentiable during the backpropagation as it is discrete. To address this issue, we estimate $f_{U}(x|s)$ using $\bm{Z}$ in equation 12. Let $\bm{e}$ denote the all-one row vector, and define $\Gamma_{Z}(\bm{T},s,x)\triangleq\text{Sigmoid}(\gamma(|\mathcal{X}|\text{Sigmoid}(\bm{Z}[s,:])-x\bm{e}))$. When $\gamma>0$ is sufficiently large, $\Gamma_{Z}(\bm{T},s,x)$ is a $(kR)$-dim vector with elements close to either 0 or $1$. Additionally, $\Gamma_{Z}(\bm{T},s,x)\bm{e}^{\top}$ approximates the number of elements in $\bm{U}[s,:]$ that are not less than $x$. We refer to $\gamma$ as the temperature parameter and estimate $f_{U}(x|s)$ for $x>0$ as follows:
$\vskip-2.0ptf_{U}(x|s)\approx\hat{f}_{U}(x|s)\triangleq\frac{1}{kR}\left[\Gamma_{Z}(\bm{T},s,x-1)\bm{e}^{\top}-\Gamma_{Z}(\bm{T},s,x)\bm{e}^{\top}\right].\vskip-2.0pt$ (13)
[…]
@@ Appendix A Notation and Definitions
$S^{+}$}, S{S} Future and current state (random variables)
@@
$S^{+}$}, S{S} Future and current state (random variables)
@@
$S^{+}$}, S{S} Future and current state (random variables)
@@ B.1 Proof of Theorem 1
$\overset{(a)}{=}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\left[H(S_{i+1}|S^{i}_{2},s_{1})-H(S_{i+1}|S^{i}_{2},U^{N},X_{i},s_{1})\right]$
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overset{(b)}{\leq}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\left[H(S_{i+1}|S_{i},s_{1})-H(S_{i+1}|S_{i},X_{i},s_{1})\right]$
@@
where (a) holds because $X_{i}$ is determined by $S^{i}$ and $U_{i}$. The equality in (b) holds if the policy is Markovian ***Correction Note: A typo appeared in the original version of this paper published at ICLR 2025. We correct it here and highlight the revision in red text. . It follows that
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overset{(a)}{\leq}\max_{\{p(u_{i}|u^{i-1})\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1})}$
$\overset{(b)}{=}\max_{\{p(x_{i}|s^{i},x^{i-1})\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1})$
$\overset{(c)}{=}\max_{\{p(x_{i}|s_{i})\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1})$ (15)
In equation B.1, the inequality in (a) holds with equality if the maximum is attained by a Markov policy, which is true as we prove below; (b) follows from the equivalence between the action-state channel and the EAS channel. We next prove (c). Let $\eta$ denote a history-dependent encoder and $\{\eta_{i}(x_{i}|s^{i},x^{i-1})\}_{i\geq 1}$ denote the associated input distribution. Furthermore, let $P^{\eta}_{i}(s,x)$ denote the probability that $S_{i}=s$ and $X_{i}=s$ conditioned on the encoder $\eta$. Note that $\eta$ can be viewed as a history-dependent policy for the MDP. Then according to MDP theory, for any history-dependent policy $\eta$, there exists a Markov policy $\eta^{\prime}$ such that $\eta$ and $\eta^{\prime}$ share the same joint probability distribution of states and actions. In particular, denote by $\eta^{\prime}_{i}(x_{i}|s_{i})$ the probability of selecting action $x_{i}$ given that the state is $s_{i}$ at time $i$. Then $\eta^{\prime}$ can be seen as a Markov encoder with input distribution $\{\eta^{\prime}_{i}(x_{i}|s_{i})\}_{i\geq 1}$. By letting
@@ C.3 Training and inference algorithm
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Parameters (k) $32.244$ $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}46.51$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}103.52$} $50.536$ $50.536$ $50.536$
FLOPs (millions) $0.3164$ $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0.4546$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.007$} $0.1989$ $0.1989$ $0.1989$
Coding Time (ms) $6.691$ $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}6.69$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}7.04$} $2.613$ $2.613$ $2.613$
Table 4: Number of parameters, FLOPs, and coding times for the Act2Comm scheme ($K=12$, $\mu=3$, $R=1/3$, $s_{in}=20$) with varying critic network training steps $s_{in}$ for both training and inference phases. Note that the backpropagation computational complexity is approximately 232–3 times that of the forward computation.